Two less than 2 times a number is 64 plus the same number. What is 3 times the number?
Answer by stanbon(75887)

Two less than 2 times a number is 64 plus the same number.
---
2x-2 = 64+x
x = 66
----
What is 3 times the number?
3*66 = 198
===============
Cheers,
Stan H.